OpenStudy (anonymous):

x=log(basea)bc , y=log(baseb)ca , z=log(basec)ab then 1+abc = ?

jhonyy9 (jhonyy9):

little help how can you rewriting x using log base a ? x=logbasea of what ?

jhonyy9 (jhonyy9):

and y rewrite using log base b and z log base c so log base a note log(a) so than x=log(a) a^x =xlog(a) a because log(a) a =1 now is it ok ?

OpenStudy (anonymous):

i am not getting anything sorry

OpenStudy (anonymous):

@jhonyy9

jhonyy9 (jhonyy9):

will resulz a^x =bc b^y =ca c^z =ab using these relations you can calcule easy 1+abc= ...

jhonyy9 (jhonyy9):

will result sorry